OpenStudy (anonymous):

Which of these is the equation for a circle with a radius of 10 and a center at (11, -2)? a.(x + 11)2 – (y – 2)2 = 100 b.(x + 11)2 + (y – 2)2 = 100 c.(x – 11)2 – (y + 2)2 = 100 d.(x – 11)2 + (y + 2)2 = 100

OpenStudy (anonymous):

i believe c

OpenStudy (anonymous):

the signs are flipped

OpenStudy (anonymous):

What do you mean signs are flipped?

OpenStudy (anonymous):

notice that 11,-2 and that corresponds to x,y .(x – 11)2 – (y + 2)2 = 100 ^ flipped ^flipped the formula is (x - h)2 + (y-k)2 = r2 ^ forgot about the plus it's D actually

OpenStudy (anonymous):

x,y is h,k

OpenStudy (anonymous):

awesome :) thank you

OpenStudy (anonymous):

(x - h)2 + (y-k)2 = r2 ^ whatever x coordinate of origin you are given here. But with flipped sign ^ same here but for y coordinate ^ here you square whatever radius they gave you.